OpenStudy (anonymous):

A particular radioactive isotope decays from 180 milligrams (mg) to 156.6 mg in 10 days. Find the half life of the isotope.

OpenStudy (anonymous):

What is the answer in days?

OpenStudy (anonymous):

The formula I am given is: \[y=y_{0}(1-r)^{x}\]

OpenStudy (dumbcow):

oh ok its same process , i just assumed it was exponential you can use either form really anyway, instead of solving for "k" first , you would solve for "r"

OpenStudy (anonymous):

Ok, so: \[156.6=180(1-r)^{10}\] ?

OpenStudy (dumbcow):

yes then \[\ln (156.6/180) = \ln (1-r)^{10}\]

OpenStudy (anonymous):

r=-0.139

OpenStudy (dumbcow):

hmm r cant be negative because 1-r must be less than one to make it a decay function

OpenStudy (dumbcow):

the exponent "10" can come out front so \[\ln (1-r) = \frac{\ln(156.6/180)}{10}\] then to cancel the log on left side take "e" raised to both sides \[\large 1-r = e^{\frac{\ln(156.6/180)}{10}} = (\frac{156.6}{180})^{1/10}\]

OpenStudy (anonymous):

I get -0.013926207, but how do I calculate the days?

OpenStudy (dumbcow):

not sure how you are getting a negative number? but to get the days, you need to solve this equation for "x" \[(1-r)^x = \frac{1}{2}\]

OpenStudy (dumbcow):

\[x = \frac{\ln (1/2)}{\ln (1-r)}\]

OpenStudy (anonymous):

ok so that is 10 days, but what is the half life?

OpenStudy (dumbcow):

what is 10 days? that is the half-life .... its the "x" value when y = 90 or y/y_0 = 1/2

OpenStudy (dumbcow):

ok originally we found that \[\ln(1-r) = \frac{\ln (156.6/180)}{10}\] plug that into \[x = \frac{\ln(1/2)}{\ln(1-r)}\] and you get your answer

OpenStudy (anonymous):

49.42547845

OpenStudy (dumbcow):

close enough ... i get 49.77 be careful with rounding...try to store answers or put entire expression in calculator

OpenStudy (anonymous):

ok yes I finally got 49.77 also...thank you for your patience...My webwork says it is correct.

OpenStudy (dumbcow):

yw